OpenStudy (anonymous):

Solve by variation of parameters: y'' + y = sinx Y'' + Y = 0 \[\lambda = \pm i\] So, \[y_{c} =Acosx+Bsinx\] How come the solutions to the complementary function are \(y_1 = cosx\) , \(y_2 = sinx\)??

OpenStudy (amistre64):

becasue e^(ix) is another way to write some trig functions

OpenStudy (anonymous):

Actually, I don't understand how we can separate them. I know when it has complex roots \(a\pm bi\) for homogeneous equation, we have the solution \[y=e^{ax}[Acos(bx)+Bsin(bx)]\], but how can we write it into two roots?

OpenStudy (amistre64):

distributive property \[y=Ae^{ax}cos(bx)+Be^{ax}sin(bx)\] voila!! 2 roots :) or rather, lets start out with the roots you found\[y_h=c_1e^{ix}+c_2e^{-ix}\]then see if we can work it into thats

OpenStudy (anonymous):

For complex roots \(a \pm bi \)\[y = c_{1} e^{(a+bi)x} + c_{2} e^{(a-bi)x}\] \[ y = e^{ax} (c_{1} e^{(bi)x} + c_{2} e^{(-bi)x})\]By \( e^{ix} = cosx + isinx \) \[ y = e^{ax} (c_{1} (cos(bx) + isin(bx) + c_{2} cos(-bx) + isin(-bx)) \]\[ y = e^{ax} (c_{1} (cos(bx) + isin(bx)) + c_{2} (cos(bx) - isin(bx))) \]\[y = e^{ax} (c_{1} cos(bx) + c_{2} cos(bx) + c_{1}isin(bx)- c_{2}isin(bx)) \]\[y = e^{ax} [(c_{1} + c_{2}) cos(bx) + (c_{1}- c_{2})isin(bx)] \]\[y = e^{ax} (A cos(bx) + Bsin(bx)) \] In this case, a = 0, b=1. So, \[y_{c} = (A cosx + Bsinx) \]

OpenStudy (anonymous):

But does that mean the two roots to the complementary function(I meant the one I have to use for finding the particular solution) are cosx and sinx?

OpenStudy (amistre64):

nice write up :) and yes, sinx and cosx are the solutions to the homogenous, some books say compliment, and are used to determine the particular solution

OpenStudy (anonymous):

So, the key idea is when I see \(y_{c} = c_1 f_{1}(x)+c_2f_{2}(x)\), then the two solutions I need for finding the particular solution are \(f_{1}(x)\) and \(f_{2}(x)\)?

OpenStudy (amistre64):

correct, are you going to use the wronskian? or go at it in the longhand mathod?

OpenStudy (amistre64):

the 2nd D of cosx = -cosx the 2nd D of sinx = -sinx y'' + y = 0 in both cases just as a double check for the homogenous part

OpenStudy (anonymous):

Wronskian? I only know I can use it to determine if the equation is linearly dependent..

OpenStudy (amistre64):

the Wronskian works like the cramer rule, or a cross product

OpenStudy (anonymous):

Oh! That's another trouble, which I'm going to post it as a new question.. I don't understand how it works!

OpenStudy (anonymous):

Should I post my problem in this post? Actually, it's the same question (equation), but I have two problems concerning this equation. The first one is (somehow) solved now.

OpenStudy (amistre64):

\[\begin{vmatrix} W_x&W&W_y\\ f_1&0&f_2\\ f'_1&g&f'_2 \end{vmatrix}\]

OpenStudy (amistre64):

\[Wx=0f'_2-gf_2\] \[Wy=gf_1-0f'_1\] \[W=f_1f'_2-f_2f'_1\]

OpenStudy (amistre64):

this is just a shortcut of the longhand version becasue the long hand ends up using a cramer rule

OpenStudy (amistre64):

let A and B be 2 unknown functions of x: A(x) + B(x) y = A(x)cos(x) + B(x)sin(x) y' =A'(x)cos(x)-A(x)sin(x) + B'(x)sin(x) +B(x)cos(x) let A'(x)cos(x) + B'(x)sin(x) represent the homogenous solution; so it equals 0 y' =-A(x)sin(x)+B(x)cos(x) y'' =-A'(x)sin(x)-A(x)cos(x)+B'(x)cos(x)-B(x)sin(x) y'' + y = sin(x) y = A(x)cos(x) +B(x)sin(x) y'' =-A'(x)sin(x)-A(x)cos(x)+B'(x)cos(x)-B(x)sin(x) ---------------------------------------------\ sin(x) =-A'(x)sin(x) +B'(x)cos(x) solve the system A' cos(x) + B'sin(x) = 0 -A' sin(x) + B'cos(x) = sin(x)

OpenStudy (anonymous):

Oh!!! Thanks! Actually, I've solved it by those two equations already. But then my teacher said something I don't understand :(

OpenStudy (amistre64):

i hope thats a reply your typing and not a lag on the site :)

OpenStudy (anonymous):

\[y'' + y =sinx\]\[Y''+Y=0\]\[Y=c_1cosx+c_2sinx\]\[y_p = u_1cosx+u_2sinx\]\[u_1' cosx+u_2'sinx=0 -(1)\]\[u_1' (-sinx)+u_2'cosx=sinx-(2)\]From (1), we get \[u_1' = -\frac{u_2'sinx}{cosx}-(3)\]Sub (3) into (2) \[( -\frac{u_2'sinx}{cosx}) (-sinx)+u_2'cosx=sinx\]\[u_2' = sinxcosx\]\[u_2 = \frac{1}{4}cos2x\] \[u_1' = -sin^2x\]\[u_1=-\frac{1}{2}\int(1-cos2x)dx=-\frac{1}{2}x +\frac{1}{4}sin2x\] \[y_p = (-\frac{1}{2}x +\frac{1}{4}sin2x)cosx+\frac{1}{4}cos2x(sinx)\]\[=-\frac{1}{2}xcosx + \frac{1}{2}(1-sin^2x)(sinx)+\frac{1}{4}(1-2sin^2x)sinx\]\[= -\frac{1}{2}xcosx+\frac{1}{4}sinx\]So, \[y=c_1cosx+c_2sinx -\frac{1}{2}xcosx+\frac{1}{4}sinx\]\[y=c_1cosx+c_2sinx -\frac{1}{2}xcosx\]

OpenStudy (anonymous):

That's how I solve this question yesterday/today. But my teacher said something confusing today. It's (somehow) related to undetermined coefficients.

OpenStudy (amistre64):

what was the confusing thing?

OpenStudy (anonymous):

Here are what I've copied in the lesson today (in latex): \[y''+y=sinx\]\[\text{Remove or destroy sinx}\]\[D=\frac{d}{dx}\]\[(D^2+1)sinx=0\]\[(D^2+1)[(D^2+1)y] = (D^2+1)sinx\] ^not quite sure of this step\[(D^2+1)^2y = 0\]\[e^{mx}(m^2+1)^2=0\]Don't understand what's going one here \( => \pm i, \pm i\) \[sinx, cosx, xcosx, xsinx\]\[y''+y=xcosx\]^Why we choose xcosx, and we can replace 0 by xcosx in the equation?\[(D^2+1)^2xcosx=0\]

OpenStudy (anonymous):

I think I got this step: \((D^2+1)[(D^2+1)y] = (D^2+1)sinx\) \[(D^2+1)sinx = 0\]Replace sinx by \((D^2+1)\) on the left and 0 by \((D^2+1)sinx\) on the right to get that step..

OpenStudy (anonymous):

*Replace sinx by \((D^2+1)y\)

OpenStudy (anonymous):

No, I think I got the concept wrong!

OpenStudy (amistre64):

i havent played around enough with using the D method to be sure whats actually going on either

OpenStudy (amistre64):

i understand it up to here \[(D^2+1)sinx=0\] but then id want to use quadratic methods to solve D^2+1=0; when D\(=\pm\sqrt{-1}\) which makes me think im not doing it correctly

OpenStudy (anonymous):

I just know the fact that \((D^2+1)sinx=0\) (sin)'' = -sinx => (sinx)''+sinx = 0 But I don't quite understand why we have to use this fact :|

OpenStudy (anonymous):

It looks like the given equation though. y'' + y = sinx => \((D^2+1)y = sinx\)

OpenStudy (amistre64):

http://www.codecogs.com/reference/maths/calculus/differential/the_d_operator.php this looks readable to me, and im looking over it to see if i can glean any insight from it

OpenStudy (anonymous):

So, replace sinx by \((D^2 +1)y\) in \((D^2+1)sinx=0\), and we get \((D^2+1)(D^2+1)y=0\)

OpenStudy (anonymous):

Nice! Now, at least I get to the step \((D^2+1)^2y=0\)

OpenStudy (anonymous):

Now, from that readable thingy, Let u = \((D^2+1)y\) So, we get \[(D^2+1)u=0\] u= \(c_1e^{ix}+c_2e^{-ix}\) ?!

OpenStudy (anonymous):

I little more explanation on what happens to sin(x)/4, please

OpenStudy (amistre64):

it just gets absorbed into the undetermined constant when we factor. \[c_2\sin(x)+\frac14\sin(x)=(c_2+\frac14)\sin(x)\] c2 + 1/4, is just another unknown constant. the solutions have to be linearly independant (unique), so if we have like terms, they are not different from each other and represent the same term.

OpenStudy (amistre64):

as opposed to the cos parts: c1 cos(x) and -1/2x cos(x) are not the same function due to the extra variable 'x'

OpenStudy (mendicant_bias):

I don't understand part of the math done where the complex root homogeneous soln general formula is derived from Euler's Formula; I'm busy right now and can't hop on this immediately, but it would be much appreciated if someone could later answer some questions fielded here.